Assets: Reserves 500, Loans 1000, Securities 500

Liabilites and Owners Equity: Deposits 1400, Debt 400

i.) Calculate (a) bank capital, (b) the leverage ratio, and (c) the percentage of assets financed by bank capital.

ii. ) By how much do assets need to depreciate for the bank’s capital to be wiped out?
Answer #1

A) Bank capital is defined as the value Or amounts invested by its investor or owners.

Bank capital = sum of Bank 's assets- sum of Bank's liabilities

Bank capital= (500+500+1400) -(400+1000)

=2400-1400

= 1000

B) :-Leverage is defined as Use of borrowed fund to supplement existing fundfund for purpose of further investment.

Most common leverage ratio is debt ratio and debt- to- equity ratio debt ratio= total debt/ total assets

Debt = 1400/2400

= 0.58 or 58 %

C):-To know the percentage of asset financed by bank is :

= total debt/ total assest

1400/2400

=0.58 Or 58%

That means 58% assets are financed bt Bank. Remaining 42 % is debt.

Sorry for second one question